OpenStudy (anonymous):

for the functins f(x)=√27x and g(x)=√3x, evaluae each expresion. (f+g)(3) (f-g)(x) (fg)(x) (f/g)(x) a (f+g)(3)=

OpenStudy (saifoo.khan):

\[\sqrt{27x} \]And\[ \sqrt{3x}\]Like that?

OpenStudy (precal):

(f/g)x \[\sqrt{27x}/\sqrt{3x}\] \[\sqrt{27/3}\] \[\sqrt{9}\] 3

OpenStudy (precal):

I just did one.